OpenStudy (anonymous):

need step by step on this one. please read for the equation. it is a factoring problem.

OpenStudy (anonymous):

\[m/m-1 + 22/m^2-1 = 13/m+1\]

Directrix (directrix):

Please add grouping symbols. By that I mean, does the problem begin as: m/(m-1) OR m/m - 1 ? m/m - 1 = 1 -1 = 0. m/(m-1) does not equal m/m - 1

OpenStudy (anonymous):

ok lets try this another way. hold on.

OpenStudy (anonymous):